OpenStudy (anonymous):

Help! How do I solve this rational equation... 1/2 - 1/ d-2 = d/ 16?

OpenStudy (anonymous):

Subtract the first two fractions,\[\frac{ 1 }{ 2 } - \frac{1}{d-2} = \frac{d-4}{2d-4}\]Put that equal to the right side\[\frac{d-4}{2d-4}=\frac{d}{16}\]Cross multiply the numbers and solve for d.

OpenStudy (anonymous):

ohhh okay thank you

OpenStudy (anonymous):

wait how'd you subtract the frac?